OpenStudy (anonymous):

A parachutist with a camera descends in free fall at a speed of 10m/s. The parachutist releases the camera at an altitude of 50m. A) How long does it take the camera to reach the ground? B) What is the velocity of the camera before it hits the ground? Here's what I did: a)V=-10m/s X=50m X=VT -50=(-10)t t=5seconds from 50m B) Final velocity=Initial velocity +gt V=(-10)+2(-9.8)(5); V=-59m/s Did I answers these questions correctly?

OpenStudy (lalaly):

The equation is: y = Vo*t + (1/2)*a*t^2 50 m = (10 m/s)t + (1/2)(9.8 m/s^2)(t^2) then use the quadratic formula to solve fr t

OpenStudy (lalaly):

final velocity = Vo + a*t = (10 m/s) + (9.8 m/s)*t

OpenStudy (anonymous):

Thank you very much!